# Point-hyperplane incidences via extremal graph theory

### v1

@@ 3 Point-hyperplane incidences with no $K_{s,s}$
For integers $d,s\geq 2$, there exists $C=C(d,s)>0$ such that the following holds for every sufficiently large $n$. Let $G=(A,B;E)$ be a bipartite graph, $|A|=n,|B|=m$ and $m=n^{\alpha}$. If $G$ contains no $K_{s,s}$ as a subgraph and does not contain the pattern $\Pi_{d}$, then
$e(G)\leq\begin{cases}Cn&\mbox{ if }\alpha\in[0,\beta_{1}],\\ Cm^{1-\frac{1}{t}}n&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ Cmn^{1-\frac{1}{d+2-t}}&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ Cm&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Let us start by addressing the cases when $\alpha\in[\alpha_{t},\beta_{t}]$ for some $t\in\{2,\dots,d\}$ or $\alpha\in[\beta_{d},\infty)$. In the latter case, we set $t:=d$. Suppose that $e(G)\geq C(m^{1-\frac{1}{t}}n+m)$ for some large constant $C$, and observe that $m^{1-\frac{1}{t}}n\geq m$ if $\alpha\in[\alpha_{t},\beta_{t}]$, and $m\geq m^{1-\frac{1}{t}}n$ if $\alpha\in[\beta_{d},\infty)$. We show that there exists a $t$-tuple of vertices $v_{1},\dots,v_{t}\in A$ which satisfies the assumptions of Lemma 3.1, which then implies that $G$ contains the pattern $\Pi_{d}$, contradiction. To do this, we use a variant of the dependent random choice method.
@@
Let us denote by $X$ the size of the common neighbourhood of $v_{1},\dots,v_{t}$, and let $Y$ be the number of $(d+1-t)$-tuples of vertices $u_{t},\dots,u_{d}\in N(v_{1},\dots,v_{t})$ which has less than $s$ common neighbours. If we show that $\mathbb{E}[X-Y]\geq d^{d+1}s$, then one can pick a good $t$-tuple $v_{1},\dots,v_{t}$ satisfying $X-Y\geq d^{d+1}s$ and delete one vertex from every $(d+1-t)$-tuple in $N(v_{1},\dotsm,v_{t})$ which has less than $s$ common neighbours. Then, one is left with the $t$-tuple $v_{1},\dots,v_{t}$ and a subset of their common neighbourhood which satisfy all conditions of Lemma 3.1.
@@
$\geq\prod_{i=1}^{t}\frac{\deg u-s-i+1}{n-i+1}>\left(\frac{\deg u-s-d}{n}\right)^{t}.$
Let $B_{0}\subseteq B$ be the set of vertices $u$ satisfying $\deg u\geq 2(d+s)$. Then, the number of edges incident to $B_{0}$ is $\sum_{u\in B_{0}}\deg u\geq C(m^{1-1/t}n+m)-2(d+s)m\geq e(G)/2$ for $C\geq 4(d+s)$. Therefore,
$\mathbb{E}[X]$ $\geq\sum_{u\in B_{0}}\left(\frac{\deg u-s-d}{n}\right)^{t}\geq\sum_{u\in B_{0}}\left(\frac{\deg u}{2n}\right)^{t}$
$\geq|B_{0}|\left(\frac{\sum_{u\in B_{0}}\deg u}{2|B_{0}|n}\right)^{t}\geq m\left(\frac{e(G)}{4mn}\right)^{t}\geq\left(\frac{e(G)}{4m^{1-1/t}n}\right)^{t},$
@@
$\mathbb{P}[u_{t},\dots,u_{d}\in N(v_{1},\dots,v_{t})]=\mathbb{P}[v_{1},\dots,v_{t}\in N(u_{t},\dots,u_{d})]\leq\left(\frac{|N(u_{t},\dots,u_{d})|}{n-s-d}\right)^{t}\leq(2s)^{t}n^{-t},$
assuming $n$ is sufficiently large with respect to $s$ and $d$. Thus, the expectation of $Y$ can be bounded as
$\mathbb{E}[Y]\leq\sum_{\begin{subarray}{c}u_{t},\dots,u_{d}\in B\\ |N(u_{t},\dots,u_{d})|<s\end{subarray}}\left(\frac{2s}{n}\right)^{t}\leq m^{d+1-t}\left(\frac{2s}{n}\right)^{t}.$
@@
$\mathbb{E}[X-Y]\geq\left(\frac{e(G)}{4m^{1-1/t}n}\right)^{t}-m^{d+1-t}\left(\frac{2s}{n}\right)^{t}\geq(C/4)^{t}-(2s)^{t}n^{\alpha(d+1-t)-t}\geq(C/4)^{t}-(2s)^{t}\geq d^{d+1}s.$
@@
$\mathbb{E}[X-Y]\geq\Big{(}\frac{Cm}{4m^{1-1/t}n}\Big{)}^{t}-m\Big{(}\frac{2s}{n}\Big{)}^{t}=\big{(}(C/4)^{d}-(2s)^{d}\big{)}\frac{m}{n^{d}}\geq(C/4)^{d}-(2s)^{d}\geq d^{d+1}s,$
@@
Observe that the rest of the cases already follow by symmetry. Since $1/\alpha_{t}=\alpha_{d+2-t}$ and $1/\beta_{t-1}=\beta_{d+2-t}$, we have that if $\alpha\in[\beta_{t-1},\alpha_{t}]$ for some $t\in\{2,\dots,d\}$, then $1/\alpha\in[\alpha_{d+2-t},\beta_{d+2-t}]$. Also, if $\alpha\in[0,\beta_{1}]$, then $1/\alpha\in[\beta_{d},\infty)$. Thus, if we reverse the roles of $m$ and $n$, together with the observation $n=m^{1/\alpha}$, we find ourselves in the setting of the cases already proved. This completes the proof of the theorem. ∎
@@
Let $d,s\geq 2$ be integers, then there exists $C=C(d,s)>0$ such that the following holds for every $\alpha>0$ and field $\mathbb{F}$. Let $\mathcal{P}$ be a set of $n$ points and $\mathcal{H}$ be a set of $m=n^{\alpha}$ hyperplanes in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free. Then
$I(\mathcal{P},\mathcal{H})\leq\begin{cases}Cn&\mbox{ if }\alpha\in[0,\beta_{1}],\\ Cm^{1-\frac{1}{t}}n&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ Cmn^{1-\frac{1}{d+2-t}}&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ Cm&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
4 Large complete bipartite subgraphs in incidence graph
The main goal of this section is to prove our lower bounds on $\operatorname{rs}_{d}(m,n,I)$ and $\operatorname{rs}_{d}(n,m,I)$ for $m\geq n$, that is, to prove Theorem 1.3. As observed before, the lower bounds are only interesting in the regime $\varepsilon>100\max\{n^{-\frac{1}{d-1}},m^{-\frac{1}{d}}\}$, otherwise taking the maximum degree vertex with its neighborhood gives a complete bipartite graph of the required size. Let us restate our theorem in this case.
@@ 4 Large complete bipartite subgraphs in incidence graph
Let $m\geq n$ be integers, $\varepsilon>0$, and $I=\varepsilon mn$. If $\varepsilon>100\max\{m^{-\frac{1}{d}},n^{-\frac{1}{d-1}}\}$, then
@@
Let us prepare the proof with some definitions. For a bipartite graph $G$, we denote by $\operatorname{rs}(G)$ the number of edges in the largest complete bipartite subgraph of $G=(A,B;E)$. Also, we say that a $t$-tuple of vertices $v_{1},\dots,v_{t}\in A$ is a good $t$-tuple if $N(v_{1},v_{2})\supsetneq N(v_{1},v_{2},v_{3})\supsetneq\cdots\supsetneq N(v_{1},\dots,v_{t})$ and $|N(v_{1},\dots,v_{t})|\geq 2$. The reason this definition is useful is the following lemma, which allows us to embed the pattern $\Pi_{d}$ using good $d$-tuples. One can think of Lemma 4.2 as a simpler version of Lemma 3.1 from Section 3.
@@
Suppose that $G$ contains a good $d$-tuple $v_{1},\dots,v_{d}$. Then $G$ contains the pattern $\Pi_{d}$.
@@
Let $d\geq 3$ be a positive integer and let $G=(A,B;E)$ be a bipartite graph with at least $\varepsilon mn$ edges, where $|A|=n$, $|B|=m$ and $n\leq m$. If $\varepsilon\geq 100\max\{n^{-\frac{1}{d-1}},m^{-\frac{1}{d}}\}$, then either $A$ contains a good $d$-tuple or
@@
Let $\gamma,\varepsilon>0$, and let $G=(A,B;E)$ be a bipartite graph such that $|A|=n$, $|B|=m$, and every vertex $u\in B$ has degree at least $\deg u\geq\frac{\varepsilon}{3}n$. If $\operatorname{rs}(G)<\frac{\varepsilon}{6}\gamma\cdot mn$, then for any set $X\subseteq B$ of size at least $\gamma m$, there exists a vertex $v\in A$ for which $|N(v)\cap X|\geq\frac{\varepsilon}{6}|X|$ and $N(v)\cap X\neq X$.
@@
Let $X\subseteq B$ be a subset of size at least $\gamma m$. The idea of this proof is to double count the edges incident to $X$. Let $V_{1}$ denote the set of vertices $v\in A$ for which $X\subseteq N(v)$ and let $V_{2}$ denote the set of vertices $v\in A$ with $|N(v)\cap X|<\frac{\varepsilon}{6}|X|$. To show the lemma, it suffices to find a vertex $v\in A$ belonging to neither $V_{1}$ nor $V_{2}$.
Suppose, for the sake of contradiction, that $V_{1}\cup V_{2}=A$. Then, $e(A,X)=e(V_{1},X)+e(V_{2},X)$. Since the degree of every vertex in $B$ is at least $\frac{\varepsilon n}{3}$, we have $e(A,X)\geq\frac{\varepsilon}{3}n|X|$. On the other hand, since the graph induced on $V_{1}\cup X$ is complete bipartite, we have $e(V_{1},X)\leq\operatorname{rs}(G)\leq\frac{\varepsilon}{6}\gamma\cdot mn$. Also, by the definition of $V_{2}$, we have $e(V_{2},X)=\sum_{v\in V_{2}}|N(v)\cap X|<|V_{2}|\cdot\frac{\varepsilon}{6}|X|\leq\frac{\varepsilon}{6}n|X|$. Hence,
$\frac{\varepsilon}{6}\gamma\cdot mn+\frac{\varepsilon}{6}n|X|>\frac{\varepsilon}{3}n|X|.$
Rearranging this inequality gives $|X|<\gamma m$, which is a contradiction to our initial assumption. ∎
@@
Let $d\geq 3$ be a positive integer and let $G=(A,B;E)$ be a bipartite graph with $\varepsilon mn$ edges, where $|A|=n$, $|B|=m$. If $\operatorname{rs}(G)\leq(\frac{\varepsilon}{6})^{d-1}mn$, then $G$ contains a good $(d-1)$-tuple $v_{1},\dots,v_{d-1}\in A$ with common neighbourhood of size at least $|N(v_{1},\dots,v_{d-1})|\geq 2\left(\frac{\varepsilon}{6}\right)^{d-1}m$.
@@
For $t=2,\dots,d-1$, we prove that there is a good $t$-tuple $v_{1},\dots,v_{t}\in A$ with $|N(v_{1},\dots,v_{t})|\geq 2\left(\frac{\varepsilon}{6}\right)^{t}m$. We prove this by induction on $t$. For $t=2$, we simply need to find a pair of vertices $v_{1},v_{2}\in A$ whose common neighbourhood has size at least $2\left(\frac{\varepsilon}{6}\right)^{2}m$. In fact, the average size of the common neighbourhood of two vertices of $A$ is $\frac{\varepsilon^{2}}{2}m$, as shown by equation (2) in Section 2, and therefore we are done.
For $t>2$, the induction hypothesis implies that $G$ contains a good $(t-1)$-tuple $v_{1},\dots,v_{t-1}\in A$ with the common neighbourhood of size $|N(v_{1},\dots,v_{t-1})|\geq 2\left(\frac{\varepsilon}{6}\right)^{t-1}m$. Let $\gamma=(\frac{\varepsilon}{6})^{t-1}$, then $\operatorname{rs}(G)\leq(\frac{\varepsilon}{6})^{d-1}mn\leq\frac{\varepsilon}{6}\gamma\cdot mn$ and $|X|\geq\gamma m$ are satisfied, so we can apply Lemma 4.4 to $X=N(v_{1},\dots,v_{t-1})$. We get that there exists a vertex $v_{t}\in A$ for which $|N(v_{1},\dots,v_{t})|\geq\frac{\varepsilon}{6}|X|\geq(\frac{\varepsilon}{6})^{t}m$ and $N(v_{1},\dots,v_{t-1})\supsetneq N(v_{1},\dots,v_{t})$. We conclude that $v_{1},\dots,v_{t}$ is a good $t$-tuple with sufficiently large common neighbourhood, completing the proof. ∎
@@
We begin the proof by showing that we can assume large minimum degree, more precisely that $\min_{v\in A}\deg v\geq\left(1-\frac{2}{d}\right)\varepsilon|B|$ and $\min_{v\in B}\deg v\geq\left(1-\frac{2}{d}\right)\varepsilon|A|$. Then, we will present the main argument, in which we strengthen Lemma 4.5 and show there exists a good $(d-1)$-tuple with common neighbourhood of size $\left(\frac{\varepsilon}{6}\right)^{d-2}m$, which allows us to apply Lemma 4.4 to find a good $d$-tuple and complete the proof.
We will consider a proof by double induction, first on $d$ and then on $m+n$. Note that if $\min\{m,n\}=1$, then the statement of the lemma is trivial. Our first step will be, as mentioned above, to show that if $G$ has a vertex of low degree, then one can remove this vertex and obtain a graph $G^{\prime}$ with $\operatorname{rs}(G^{\prime})\geq(\frac{\varepsilon}{100})^{d-1}mn$.
@@
Since $A$ and $B$ are symmetric, we will only discuss the case when $G$ has a vertex $v\in A$ with degree $\deg v\leq\left(1-\frac{2}{d}\right)\varepsilon|B|$. Define the graph $G^{\prime}=G-\{v\}$, which has parts of size $n^{\prime}=n-1$ and $m^{\prime}=m$. Furthermore, $G^{\prime}$ has $\varepsilon^{\prime}m^{\prime}n^{\prime}$ edges, where $\varepsilon^{\prime}m^{\prime}n^{\prime}\geq\varepsilon mn-\left(1-\frac{2}{d}\right)\varepsilon m$. Thus, if $G-\{v\}$ does not contain the pattern $\Pi_{d}$, we conclude
$\operatorname{rs}(G)$ $\geq\operatorname{rs}(G^{\prime})\geq\left(\frac{\varepsilon^{\prime}}{100}\right)^{d-1}m^{\prime}n^{\prime}\geq\left(\frac{\varepsilon mn-\left(1-\frac{2}{d}\right)\varepsilon m}{100(n-1)m}\right)^{d-1}(n-1)m$
$=\left(\Big{(}\frac{\varepsilon}{100}\Big{)}^{d-1}mn\right)\cdot\frac{n-1}{n}\left(\frac{n-(1-\frac{2}{d})}{n-1}\right)^{d-1}\geq\Big{(}\frac{\varepsilon}{100}\Big{)}^{d-1}mn.$

### v2

@@ 3.1 Upper bounds for $K_{s,s}$-free incidence graphs
For integers $d,s\geq 2$, there exists $C=C(d,s)>0$ such that the following holds for every sufficiently large $m$. Let $G=(A,B;E)$ be a bipartite graph, $|A|=m,|B|=n$ and $n=m^{\alpha}$. If $G$ contains no $K_{s,s}$ as a subgraph and does not contain the pattern $\Pi_{d}$, then
$e(G)\leq\begin{cases}Cm&\mbox{ if }\alpha\in[0,\beta_{1}],\\ Cmn^{1-\frac{1}{t}}&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ Cm^{1-\frac{1}{d+2-t}}n&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ Cn&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Let us start by addressing the cases when $\alpha\in[\alpha_{t},\beta_{t}]$ for some $t\in\{2,\dots,d\}$ or $\alpha\in[\beta_{d},\infty)$. In the latter case, we set $t:=d$. Suppose that $e(G)\geq C(mn^{1-\frac{1}{t}}+n)$ for some large constant $C$, and observe that $mn^{1-\frac{1}{t}}\geq n$ if $\alpha\in[\alpha_{t},\beta_{t}]$, and $n\geq mn^{1-\frac{1}{t}}$ if $\alpha\in[\beta_{d},\infty)$. We show that there exists a $t$-tuple of vertices $v_{1},\dots,v_{t}\in A$ which satisfies the assumptions of Lemma 3.1, which then implies that $G$ contains the pattern $\Pi_{d}$, contradiction. To do this, we use a variant of the dependent random choice method.
@@
Let us denote by $X$ the size of the common neighbourhood of $v_{1},\dots,v_{t}$, and let $Y$ be the number of $(d+1-t)$-tuples of vertices $u_{t},\dots,u_{d}\in N(v_{1},\dots,v_{t})$ which has less than $s$ common neighbours. If we show that $\mathbb{E}[X-Y]\geq d^{d+1}s$, then one can pick a good $t$-tuple $v_{1},\dots,v_{t}$ satisfying $X-Y\geq d^{d+1}s$ and delete one vertex from every $(d+1-t)$-tuple in $N(v_{1},\dots,v_{t})$ which has less than $s$ common neighbours. Then, one is left with the $t$-tuple $v_{1},\dots,v_{t}$ and a subset of their common neighbourhood which satisfy all conditions of Lemma 3.1.
@@
$\geq\prod_{i=1}^{t}\frac{\deg u-s-i+1}{m-i+1}>\left(\frac{\deg u-s-d}{m}\right)^{t}.$
Let $B_{0}\subseteq B$ be the set of vertices $u$ satisfying $\deg u\geq 2(d+s)$. Then, the number of edges incident to $B_{0}$ is $\sum_{u\in B_{0}}\deg u\geq C(mn^{1-1/t}+n)-2(d+s)n\geq e(G)/2$ for $C\geq 4(d+s)$. Therefore,
$\mathbb{E}[X]$ $\geq\sum_{u\in B_{0}}\left(\frac{\deg u-s-d}{m}\right)^{t}\geq\sum_{u\in B_{0}}\left(\frac{\deg u}{2m}\right)^{t}$
$\geq|B_{0}|\left(\frac{\sum_{u\in B_{0}}\deg u}{2|B_{0}|m}\right)^{t}\geq n\left(\frac{e(G)}{4mn}\right)^{t}\geq\left(\frac{e(G)}{4mn^{1-1/t}}\right)^{t},$
@@
$\mathbb{P}[u_{t},\dots,u_{d}\in N(v_{1},\dots,v_{t})]=\mathbb{P}[v_{1},\dots,v_{t}\in N(u_{t},\dots,u_{d})]\leq\left(\frac{|N(u_{t},\dots,u_{d})|}{m-s-d}\right)^{t}\leq(2s)^{t}m^{-t},$
assuming $m$ is sufficiently large with respect to $s$ and $d$. Thus, the expectation of $Y$ can be bounded as
$\mathbb{E}[Y]\leq\sum_{\begin{subarray}{c}u_{t},\dots,u_{d}\in B\\ |N(u_{t},\dots,u_{d})|<s\end{subarray}}\left(\frac{2s}{m}\right)^{t}\leq n^{d+1-t}\left(\frac{2s}{m}\right)^{t}.$
@@
$\mathbb{E}[X-Y]\geq\left(\frac{e(G)}{4mn^{1-1/t}}\right)^{t}-n^{d+1-t}\left(\frac{2s}{m}\right)^{t}\geq(C/4)^{t}-(2s)^{t}m^{\alpha(d+1-t)-t}\geq(C/4)^{t}-(2s)^{t}\geq d^{d+1}s.$
@@
$\mathbb{E}[X-Y]\geq\Big{(}\frac{Cn}{4mn^{1-1/t}}\Big{)}^{t}-m\Big{(}\frac{2s}{m}\Big{)}^{t}=\big{(}(C/4)^{d}-(2s)^{d}\big{)}\frac{n}{m^{d}}\geq(C/4)^{d}-(2s)^{d}\geq d^{d+1}s,$
@@
Observe that the rest of the cases already follow by symmetry. Since $1/\alpha_{t}=\alpha_{d+2-t}$ and $1/\beta_{t-1}=\beta_{d+2-t}$, we have that if $\alpha\in[\beta_{t-1},\alpha_{t}]$ for some $t\in\{2,\dots,d\}$, then $1/\alpha\in[\alpha_{d+2-t},\beta_{d+2-t}]$. Also, if $\alpha\in[0,\beta_{1}]$, then $1/\alpha\in[\beta_{d},\infty)$. Thus, if we reverse the roles of $m$ and $n$, together with the observation $m=n^{1/\alpha}$, we find ourselves in the setting of the cases already proved. This completes the proof of the theorem. ∎
@@
Let $d,s\geq 2$ be integers, then there exists $C=C(d,s)>0$ such that the following holds for every $\alpha>0$ and field $\mathbb{F}$. Let $\mathcal{P}$ be a set of $m$ points and $\mathcal{H}$ be a set of $n=m^{\alpha}$ hyperplanes in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free. Then
$I(\mathcal{P},\mathcal{H})\leq\begin{cases}Cm&\mbox{ if }\alpha\in[0,\beta_{1}],\\ Cmn^{1-\frac{1}{t}}&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ Cm^{1-\frac{1}{d+2-t}}n&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ Cn&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
3.2 Large complete bipartite subgraph of dense incidence graphs
The main goal of this section is to prove our lower bounds on $\operatorname{rs}_{d}(m,n,I)$ and $\operatorname{rs}_{d}(n,m,I)$ for $m\leq n$, that is, to prove Theorem 1.3. Recall that $\operatorname{rs}_{d}(n,m,I)=\min\operatorname{rs}(\mathcal{P},\mathcal{H})$, where the minimum is taken over all fields $\mathbb{F}$, all sets $\mathcal{P}\subset\mathbb{F}^{d}$ of at most $m$ points, all sets $\mathcal{H}$ of at most $n$ hyperplanes in $\mathbb{F}^{d}$, that satisfy $I(\mathcal{P},\mathcal{H})\geq I$. As observed before, the lower bounds are only interesting in the regime $\varepsilon>100\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}$, otherwise taking the maximum degree vertex with its neighborhood gives a complete bipartite graph of the required size. Let us restate our theorem in this case.
@@ 3.2 Large complete bipartite subgraph of dense incidence graphs
Let $m\leq n$ be integers, $\varepsilon>0$, and $I=\varepsilon mn$. If $\varepsilon>100\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}$, then
@@
Let us prepare the proof with some definitions. For a bipartite graph $G$, we denote by $\operatorname{rs}(G)$ the number of edges in the largest complete bipartite subgraph of $G=(A,B;E)$. Also, we say that a $t$-tuple of vertices $(v_{1},\dots,v_{t})\in A^{t}$ is a good $t$-tuple if $N(v_{1},v_{2})\supsetneq N(v_{1},v_{2},v_{3})\supsetneq\cdots\supsetneq N(v_{1},\dots,v_{t})$ and $|N(v_{1},\dots,v_{t})|\geq 2$. The reason this definition is useful is the following lemma, which allows us to embed the pattern $\Pi_{d}$ using good $d$-tuples. One can think of Lemma 3.5 as a simpler version of Lemma 3.1 from Section 3.1.
@@
Suppose that $G$ contains a good $d$-tuple $(v_{1},\dots,v_{d})$. Then $G$ contains the pattern $\Pi_{d}$.
@@
Let $d\geq 2$ be a positive integer and let $G=(A,B;E)$ be a bipartite graph with at least $\varepsilon mn$ edges, where $|A|=m$, $|B|=n$ and $n\geq m$. If $\varepsilon\geq 100\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}$, then either $A$ contains a good $d$-tuple or
@@
Let $\gamma,\varepsilon>0$, and let $G=(A,B;E)$ be a bipartite graph such that $|A|=m$, $|B|=n$, and every vertex $u\in B$ has degree at least $\deg u\geq\frac{\varepsilon}{3}n$. If $\operatorname{rs}(G)<\frac{\varepsilon}{6}\gamma\cdot mn$, then for any set $X\subseteq B$ of size at least $\gamma n$, there exists a vertex $v\in A$ for which $|N(v)\cap X|\geq\frac{\varepsilon}{6}|X|$ and $N(v)\cap X\neq X$.
@@
Let $X\subseteq B$ be a subset of size at least $\gamma n$. The idea of this proof is to double count the edges incident to $X$. Let $V_{1}$ denote the set of vertices $v\in A$ for which $X\subseteq N(v)$ and let $V_{2}$ denote the set of vertices $v\in A$ with $|N(v)\cap X|<\frac{\varepsilon}{6}|X|$. To show the lemma, it suffices to find a vertex $v\in A$ belonging to neither $V_{1}$ nor $V_{2}$.
Suppose, for the sake of contradiction, that $V_{1}\cup V_{2}=A$. Then, $e(A,X)=e(V_{1},X)+e(V_{2},X)$. Since the degree of every vertex in $B$ is at least $\frac{\varepsilon m}{3}$, we have $e(A,X)\geq\frac{\varepsilon}{3}m|X|$. On the other hand, since the graph induced on $V_{1}\cup X$ is complete bipartite, we have $e(V_{1},X)\leq\operatorname{rs}(G)\leq\frac{\varepsilon}{6}\gamma\cdot mn$. Also, by the definition of $V_{2}$, we have $e(V_{2},X)=\sum_{v\in V_{2}}|N(v)\cap X|<|V_{2}|\cdot\frac{\varepsilon}{6}|X|\leq\frac{\varepsilon}{6}m|X|$. Hence,
$\frac{\varepsilon}{6}\gamma\cdot mn+\frac{\varepsilon}{6}m|X|>\frac{\varepsilon}{3}m|X|.$
Rearranging this inequality gives $|X|<\gamma n$, which is a contradiction to our initial assumption. ∎
@@
Let $d\geq 3$ be a positive integer and let $G=(A,B;E)$ be a bipartite graph with $\varepsilon mn$ edges, where $|A|=m$, $|B|=n$. If $\operatorname{rs}(G)\leq(\frac{\varepsilon}{6})^{d-1}mn$, then $G$ contains a good $(d-1)$-tuple $(v_{1},\dots,v_{d-1})\in A^{d-1}$ with common neighbourhood of size at least $|N(v_{1},\dots,v_{d-1})|\geq 2\left(\frac{\varepsilon}{6}\right)^{d-1}n$.
@@
For $t=2,\dots,d-1$, we prove that there is a good $t$-tuple $(v_{1},\dots,v_{t})\in A^{t}$ with $|N(v_{1},\dots,v_{t})|\geq 2\left(\frac{\varepsilon}{6}\right)^{t}n$. We prove this by induction on $t$. For $t=2$, we simply need to find a pair of vertices $v_{1},v_{2}\in A$ whose common neighbourhood has size at least $2\left(\frac{\varepsilon}{6}\right)^{2}n$. Since the average degree of vertices in $B$ is at least $\varepsilon m$, the average size of the common neighbourhood of two vertices of $A$ is $n\binom{\varepsilon m}{2}/\binom{m}{2}\geq\frac{\varepsilon^{2}}{2}n$, which is sufficient.
For $t>2$, the induction hypothesis implies that $G$ contains a good $(t-1)$-tuple $(v_{1},\dots,v_{t-1})\in A^{t-1}$ with the common neighbourhood of size $|N(v_{1},\dots,v_{t-1})|\geq 2\left(\frac{\varepsilon}{6}\right)^{t-1}n$. Let $\gamma=(\frac{\varepsilon}{6})^{t-1}$, then $\operatorname{rs}(G)\leq(\frac{\varepsilon}{6})^{d-1}mn\leq\frac{\varepsilon}{6}\gamma\cdot mn$ and $|X|\geq\gamma n$ are satisfied, so we can apply Lemma 3.7 to $X=N(v_{1},\dots,v_{t-1})$. We get that there exists a vertex $v_{t}\in A$ for which $|N(v_{1},\dots,v_{t})|\geq\frac{\varepsilon}{6}|X|\geq(\frac{\varepsilon}{6})^{t}n$ and $N(v_{1},\dots,v_{t-1})\supsetneq N(v_{1},\dots,v_{t})$. We conclude that $(v_{1},\dots,v_{t})$ is a good $t$-tuple with sufficiently large common neighbourhood, completing the proof. ∎
@@
Assume that the statement is false, and let $G$ be a counterexample with $d$ minimal, and among those $\min\{m,n\}$ is also minimal. That is, $G$ has at least $\varepsilon mn$ edges, $G$ contains no good $d$-tuple and $\operatorname{rs}(G)<(\frac{\varepsilon}{100})^{d-1}mn$. Clearly, we have $\min\{m,n\}\geq 2$. Also, if $d=2$, finding a good $d$-tuple corresponds to finding a pair of vertices of $A$ with at least two common neighbours, i.e. finding a cycle of length $4$ in $G$. Since $|E(G)|\geq\varepsilon mn\geq 100\max\{n^{-1/2},m^{-1}\}mn\geq mn^{1/2}+n$, a simple application of the classical Kővári-Sós-Turán theorem shows that $G$ must have a cycle of length $4$, which suffices. Hence, we may assume that $d\geq 3$.
@@
First, we show that $\min_{v\in A}\deg v\geq\left(1-\frac{2}{d}\right)\varepsilon|B|$ and $\min_{v\in B}\deg v\geq\left(1-\frac{2}{d}\right)\varepsilon|A|$. Since $A$ and $B$ are symmetric, we only discuss the case when $G$ has a vertex $v\in A$ with degree $\deg v\leq\left(1-\frac{2}{d}\right)\varepsilon|B|$. Define the graph $G^{\prime}=G-\{v\}$, which has parts of size $m^{\prime}=m-1$ and $n^{\prime}=n$. Furthermore, $G^{\prime}$ has $\varepsilon^{\prime}m^{\prime}n^{\prime}$ edges, where $\varepsilon^{\prime}m^{\prime}n^{\prime}\geq\varepsilon mn-\left(1-\frac{2}{d}\right)\varepsilon n$. Thus, as $G^{\prime}$ does not contain a good $d$-tuple and by the minimality of $G$, we conclude
$\operatorname{rs}(G)$ $\geq\operatorname{rs}(G^{\prime})\geq\left(\frac{\varepsilon^{\prime}}{100}\right)^{d-1}m^{\prime}n^{\prime}\geq\left(\frac{\varepsilon mn-\left(1-\frac{2}{d}\right)\varepsilon n}{100(m-1)n}\right)^{d-1}(m-1)n$
$=\left(\Big{(}\frac{\varepsilon}{100}\Big{)}^{d-1}mn\right)\cdot\frac{m-1}{m}\left(\frac{m-(1-\frac{2}{d})}{m-1}\right)^{d-1}\geq\Big{(}\frac{\varepsilon}{100}\Big{)}^{d-1}mn.$